There are four children in one family. Adam is 3 years older than Bob, who is 2 times as old as Claire. David is 1 year younger than Claire. If their combined age is 20 years, find the age of each child.

Answer by rfer(16322)

x=clare=3
x-1=david=2
2x=bob=6
2x+3=adam=9
-------------
x+x-1+2x+2x+3=20
6x=20-2
x=18/6
x=3
x-1=2
2x=6
2x+3=9